OpenStudy (anonymous):

Given log2=0.3010 and log3=0.4771 (all base 10), find log(72)^1/3.

OpenStudy (anonymous):

Use Law of Logs, log(72)^(1/3) = 1/3 log 72 Now, log 72 = log(3^2 x 2^3)= log (3^2) + log(2^3) = 2 log 3 + 3 log 2 So your answer would be 1/3(2log 3 + 3 log 2) Use log 2 and log 3 from the stem of the question and you'll have your answer.